OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18 Please show all of your work

OpenStudy (anonymous):

the slope of the equation 7x-10y=18 is -10y=-7x+18 y=(7/10)x-18/7 so slope is 7/10 then because its perpendicular then the slope of the other line is -10/7 so slope intercept equation is y=(-10/7)x+b...... then solve for b by plugging in the point (14,-2) for x and y -2=-20+b.... so b=18 and answer is y=(-10/7)x+18

OpenStudy (anonymous):

y=7/10x-18/7 7/10 = slope 7/10*x=-1 x=-10/7 -10/7(x,y)(14,-2) change in y ------- change in x y+2 = -10/7 ---- x-14 7(y+2)=-10(x-14) 7y+14=-10x+140 y=-10/7x+20

OpenStudy (anonymous):

it should be +18 not 20

OpenStudy (anonymous):

y=-10/7x+18